OpenStudy (anonymous):

The number of bacteria N in a culture is given by the model N = 241e^(kt) where t is the time in hours. If N = 423 when t = 18, estimate the time required for the population to double in size. Answer in units of hours

OpenStudy (anonymous):

Do you know how to find k?

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

plug the numbers in ... \[423 = 241e ^{18k}\]

OpenStudy (anonymous):

divide both sides by 241

OpenStudy (anonymous):

423/241 = e^(18k)

OpenStudy (anonymous):

can you find k from here ?

OpenStudy (anonymous):

take the natural log (ln) of both sides ...can you do this?

OpenStudy (anonymous):

so k=ln 1.75518672 ?

OpenStudy (anonymous):

er

OpenStudy (anonymous):

divide by 18

OpenStudy (anonymous):

ln 1.75518672/18=k

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so then how do we find neccesary to double pop?

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

erm

OpenStudy (anonymous):

wait a sec ...

OpenStudy (anonymous):

ln (846/241) /ln 0.03125=T

OpenStudy (anonymous):

but i got a negative number >.<

OpenStudy (anonymous):

does it sound reasonable to you?

OpenStudy (anonymous):

idk.. probably.....

OpenStudy (anonymous):

I used Wolframalpha.com

OpenStudy (anonymous):

hmm according to my thingy its nt right >.<

OpenStudy (anonymous):

do you have the answer?

OpenStudy (anonymous):

no..i just know its wrong..its technically an online assignment tht checks whether or not it is right and it isnt right..

OpenStudy (anonymous):

I'll do it by hand ... hold on

OpenStudy (anonymous):

maybe because I round it ... to the nearest whole number ?

OpenStudy (anonymous):

usually its to nearest 3 decimal places i think

OpenStudy (anonymous):

that's the problem with internet math ... what system are you using?

OpenStudy (anonymous):

system?

OpenStudy (anonymous):

how many trials are you allowed?

OpenStudy (anonymous):

Bb, WebAssign , etc ...

OpenStudy (anonymous):

well severeal poinnts are taken away withc each trial so prolly 1-3

OpenStudy (anonymous):

did you copy the problem exactly ?

OpenStudy (anonymous):

yea i can give u the link if u want lol

OpenStudy (anonymous):

let me see the link

OpenStudy (anonymous):

actually i cant really get the link..

OpenStudy (anonymous):

never mind ... I think I know what I did wrong ...

OpenStudy (anonymous):

the population at t =0 is the initial pop not 423 ...

OpenStudy (anonymous):

initial pop No = 241

OpenStudy (anonymous):

double 241 = 482 ... use 482 instead of 846

OpenStudy (anonymous):

check and let me know ...it should be ok !

OpenStudy (anonymous):

its nt correct >.< why do i always get these hard problems ...

OpenStudy (anonymous):

darn ...I need a TI ... the answer doesn't make sense ... look at this : 18 hrs pop grew from 241 to 423 .. how long does it take to get to 482 ? not that long based on the growth rate ...right ? do you see why 64 hrs doesn't make sense ?

OpenStudy (anonymous):

yea my teacher had made all these questions is demented and is an epic troll checking every little mistake taking off points and makes something really easy always difficult ...well thx anyways for ur effort man i appreciate it

OpenStudy (anonymous):

ln(2)/0.0312541803086457207858349902725782990652095474249708

OpenStudy (anonymous):

I forgot to take the log of the right side ...darn it ! I'm tired ...sorry !

OpenStudy (anonymous):

22.1777430639639182994734144027519640627816489436816 hrs ..it is good for sure !

OpenStudy (anonymous):

22.17774

OpenStudy (anonymous):

Y/N ...so I can go to sleep ... +_+